OpenStudy (anonymous):

Write an equation of a circle with the given center and radius.center (4, 5), radius 7

OpenStudy (anonymous):

these are my options:

OpenStudy (ikram002p):

do u know how to write a circle equation ?

OpenStudy (ikram002p):

ok , for a center\( (x_0, y_0)\) and radius \(r\) the circle equation is \(\Huge \color{blue}{(x-x_0)^2+(y-y_0)^2=r^2}\)

OpenStudy (ikram002p):

so for ur center and radius could you find it ?

OpenStudy (ikram002p):

so what you got ?

OpenStudy (anonymous):

i got C

OpenStudy (ikram002p):

humm no put 4 instead of \(x_0\) put 5 instead of \(y_0\) and 7 instead of r try it

OpenStudy (anonymous):

ooh it would be A..

OpenStudy (ikram002p):

np :( dnt forget the equation must be r^2 so it should be 7^2 =49 not 7 okey ?

OpenStudy (anonymous):

oohh...so it would be B..:)

OpenStudy (ikram002p):

yess uve got it :D